### v2

# ATLAS: Automated Tree-based Language Analysis System for C and C++ Source Programs

Abstract.
Analyzing non-compilable C/C++ submodules without a resolved build environment remains a critical bottleneck for industrial software evolution. Traditional static analysis tools often fail in these scenarios due to their reliance on successful compilation, while Large Language Models (LLMs) lack the structural context necessary to reason about complex program logic. We introduce ATLAS, a Python-based Command-Line Interface (CLI). It is characterized by: (i) inter-procedural, type-aware analysis that captures dependencies across function boundaries; (ii) support for both full-project and partial analysis of non-compilable projects; (iii) graph optimizations, such as variable collapsing and node blacklisting; and (iv) synchronized multi-view graphs that align syntax, execution paths, and data-flow logic. Evaluating ATLAS with DeepSeek V3.2 for automated test generation demonstrates a 34.71% increase in line coverage and 32.66% in branch coverage, achieving performance on par with and even exceeding the industry-standard symbolic execution tool KLEE on complex projects. With polynomial scalability, ATLAS provides a robust infrastructure for generating the information-dense datasets required by next-generation, graph-aware Machine Learning for Software Engineering (ML4SE) models.
Video demonstration: https://youtu.be/xEJDNYkiD4s
Tool github repository: https://github.com/jaid-monwar/ATLAS-multi-view-code-representation-tool.git
Static Program Analysis, Control Flow, Data Flow Analysis, C/C++
conference: ; ;
1. Introduction
Programming languages like C and C++ are considered still as the foundation of performance-critical software, such as operating systems and embedded firmware (Stroustrup, 2020), where hardware proximity and resource efficiency are paramount. However, as C/C++ based systems scale, managing their evolution becomes increasingly complex (Brooks, 1987), particularly when analyzing sub-modules that are non-compilable or contain unresolved dependencies. Consequently, the field has transitioned toward the Machine Learning for Software Engineering (ML4SE) to automate tasks like test generation (Fontes and Gay, 2023) (Durelli et al., 2019), code translation (Tufano et al., 2019) (Ibrahimzada et al., 2025), source code classification (Mou et al., 2016), and code summarization (Haiduc et al., 2010) (Jiang et al., 2017). Within this paradigm, Large Language Models (LLMs) offer a flexible alternative to traditional tools. These can derive semantic context even from unbuildable projects by processing code as structured textual sequences (Wang et al., 2025) (Panichella et al., 2013) (Li et al., 2025).
However, LLMs are limited by their probabilistic nature, often prioritizing token adjacency over the deep, non-linear relationships that dictate program behavior (Anand et al., 2024) (Pendyala and Thakur, 2025). Bridging this ’semantic gap’ requires effective representation learning. Such techniques are essential for training and evaluating ML models while also enhancing the reasoning capabilities of LLMs. Empirical evidence (Liu and Jabbarvand, 2025) (Samoaa et al., 2022) shows that performance improves significantly when models utilize multi-view data that explicitly encodes these underlying logical relationships.
To bridge this gap, research has shifted toward representing code as integrated graphs (Chimalakonda et al., 2023). The Abstract Syntax Tree (AST) encodes the hierarchical syntactic structure of code constructs, the Control Flow Graph (CFG) models execution paths, and the Data Flow Graph (DFG) captures the variable dependencies. However, their integration into a multi-view representation allows models to holistically capture program logic. This approach is validated by CodeSAM (Mathai et al., 2024), which uses attention-masking to improve code search, and PANTA (Gu et al., 2025), which leverages CFG analysis to increase the test coverage by 23%. However, these successes are largely centered on Java, and C/C++ projects lack such high-level abstractions and require significantly more code to implement equivalent logic, exhausting LLM context windows more rapidly(Chimalakonda et al., 2023). Moreover, this difficulty is further complicated by C-specific hurdles like pointer aliasing, and type obscuration, which remain under-addressed (Schubert et al., 2021) in representation learning research.
To address these challenges, this paper introduces ATLAS, a Python-based Command-Line Interface (CLI) designed to function as a high-performance data generation engine for ML4SE. ATLAS automates the extraction of correct, aligned multi-view representations from large-scale C/C++ repositories. By resolving inter-procedural dependencies and offering graph-optimization options, ATLAS allows developers and researchers to quickly generate information-dense datasets that are directly ingestible by modern Machine Learning (ML) and Artificial Intelligence (AI) frameworks. The key contributions of this work are:
  • Statement-level, inter-procedural CFG and type-aware DFG generation.
  • Support for multi-file analysis of both compilable and non-compilable C/C++ projects.
  • Graph optimizations, including variable collapsing (merging redundant references) and node blacklisting (filtering irrelevant types).
  • Unified multi-view representations synchronizing syntax, control, and data flow for simultaneous learning.
2. Related Work
Compiler infrastructures like LLVM and GCC provide high-fidelity analysis but require fully resolvable build environments (Dubniczky et al., 2025), making them impractical for fragmented codebases (e.g., forcing full compilation just to analyze a small submodule). Furthermore, Intermediate Representation (IR)-based analysis inherently discards source-level semantics, limiting the efficacy of downstream models (Lattner et al., 2020). In contrast, ATLAS operates on source ASTs, preserving semantic metadata for non-compilable projects and natively modeling constructs like templates and lambdas.
Additionally, static tools like Joern (14) provide multi-language code-views but lack a native mechanism to customize or combine these representations into a unified structure. This often results in isolated, function-level graphs that cannot resolve the long-range dependencies required for a comprehensive project-level analysis. Similarly, while Tree-sitter (Brunsfeld and Contributors, 2023) parses syntax, it cannot resolve execution logic or data-flow. In contrast, ATLAS employs custom inter-procedural algorithms to precisely model pointer indirection and class-member state at the statement level. While COMEX (Das et al., 2023) targets memory-safe Java and C#, ATLAS addresses unmanaged C/C++ semantics through specialized, type-aware analysis.
ATLAS further addresses two critical gaps in the current landscape. First, it provides statement-level DFGs for C and C++, whereas existing research tools like PhASAR (Schubert et al., 2019) focus primarily on IR. Second, it resolves the fragmentation of current toolchains, where developers must currently stitch together disparate tools, such as Clang for ASTs and Joern (14) for CFGs, using custom scripts. This fragmented process is inefficient and lacks cross-view alignment. ATLAS provides a unified framework for aligned multi-view graphs, eliminating redundant analysis and streamlining ML4SE dataset pipelines.
3. ATLAS Tool
Figure 1. Architecture of ATLAS
Figure 2. Visual representations of code analysis structures by ATLAS: (a) AST showing the hierarchical syntactic structure of the source code. (b) CFG illustrating the execution paths and control logic between statements. (c) DFG depicting the flow of data values and dependencies between variables and operations.
ATLAS follows a two-step workflow (Fig. 1). At first, the Preprocessor and Parser normalize multi-file projects into a Parsed Code Model enriched with scope, type, and declaration-use metadata. Second, the Combined Code View Generator constructs aligned graphs via four modules: the AST, CFG, and DFG Drivers, and the Unified Multi-view Graph. These components utilize Reaching Definition Analysis (RDA) and inter-procedural mapping to ensure consistent node alignment across all views. Finally, ATLAS exports these representations in JSON, DOT, or PNG, linking a program’s syntactic structure directly to its execution behavior.
3.1. Preprocessor and Parser
The Preprocessor and Parser generate a unified Parsed Code Model. The Preprocessor normalizes the codebase by consolidating multi-file projects, removing non-semantic elements (e.g., comments and empty lines), and resolving global macro and type definitions. Subsequently, the Parser leverages Tree-sitter to generate concrete syntax trees (CSTs), which are traversed to build an enriched symbol table mapping scope hierarchies, data types, and declaration-use relationships. This metadata layer serves as the foundational reference for all subsequent graph generation.
3.2. Code Representation Generation
The Combined Code View Generator utilizes the enriched symbol table to construct aligned AST, CFG, and DFG views. This shared reference ensures node-level consistency for unified multi-view analysis. Construction logic and refinements for each representation are as follows.
3.2.1. AST
The CST produced in the preprocessing step preserves all syntactic details, which ATLAS traverses to build an AST index mapping each node’s position and type to a unique identifier. To refine this representation, ATLAS prunes non-semantic tokens, such as brackets and semicolons, to produce a core AST, as shown in Fig. 2(a). The AST Driver further provides two customization mechanisms to control graph granularity: (i) Variable Collapsing, which unifies all occurrences of a specific variable identifier into a single node to reduce redundancy, and (ii) Node Blacklisting, which allows users to omit irrelevant node categories (e.g., comment, modifier) and their children. These refinements allow users to distill the program into its core logical elements, effectively mitigating the computational overhead and overfitting risks associated with high-dimensional syntax.
3.2.2. CFG
The CFG Driver generates statement-level, inter-procedural representations where each node corresponds to a source statement and edges map all potential execution transitions. Unlike traditional tools that produce isolated function-level graphs, ATLAS establishes global connectivity by resolving function calls as directed edges from caller statements to callee definitions. This capability is demonstrated in Fig. 2(b), where the ternary operator on line 20 triggers indirect calls to f1() and f2() via function pointers; ATLAS correctly models this ambiguity by generating branching edges to both function definitions. Furthermore, the tool natively supports recursion by representing it as self-referential edges, as seen in the recursive call within f1() on line 7. By unifying these complex control structures into a single connected graph, ATLAS enables a comprehensive trace of program execution often lost in intra-procedural analysis.
3.2.3. DFG
The DFG Driver employs Reaching Definition Analysis (RDA) to construct a directed graph where nodes represent variable definitions, uses, and modifications. RDA ensures that data dependencies reflect actual program execution semantics, rather than mere textual proximity, by tracking which variable definitions reach a program point without being overwritten. This enables inter-procedural, type-aware tracking as demonstrated in Fig. 2(c). ATLAS accurately models pass-by-reference semantics by linking variables to their aliased parameters across function boundaries. For instance, the definition of k on Line 16 is traced through its reference as parameter a in f1(). Furthermore, the tool tracks class-member state throughout an object’s lifecycle. It connects the initialization of member variable x in the constructor (Line 7) to its subsequent usage within the member function f1() on Line 11.
3.2.4. Combined Multi-View Graph
ATLAS merges these complementary views into a unified multi-view graph by preserving consistent node identities across all representations. This alignment allows downstream ML/DL models to learn from syntactic, control-flow, and data-flow properties simultaneously. The tool supports fifteen combination variants, including pairwise and triple-view integrations, all of which remain compatible with user-defined collapsing and blacklisting. By assigning distinct type attributes to each edge class, ATLAS enables models to selectively prioritize specific program semantics over detail-heavy syntax.
3.3. Visualizer
The Visualizer manages the exportation of aligned graphs into three different formats. While the JSON output is optimized for Machine Learning (ML) integration, the DOT and PNG formats facilitate human-centric debugging and structural inspection. For multi-view representations, the component employs color-coded edge classes to distinguish between syntactic, control, and data-flow properties. Additionally, the Visualizer implements specialized C++ formatting, such as escaping namespace qualifiers and reserved keywords, to ensure high-fidelity rendering of complex codebase structures.
4. ATLAS In Practice
Table 1. Coverage comparison across DeepSeek with CFG, DeepSeek without CFG, and KLEE on TheAlgorithms C dataset
Identifier #Function #Line #Branch Function Cov. (%) Line Cov. (%) Branch Cov. (%)
DS w/ CFG DS w/o CFG KLEE DS w/ CFG DS w/o CFG KLEE DS w/ CFG DS w/o CFG KLEE
red_black_tree 11 309 206 100.00 70.00 100.00 90.00 21.70 82.00 91.30 15.00 43.80
multikey_quick_sort 27 301 196 100.00 41.20 100.00 86.40 34.50 77.60 92.90 39.00 65.80
hangman 14 170 94 100.00 66.70 70.60 92.40 53.10 64.50 87.20 46.20 56.30
min_heap 17 166 122 100.00 100.00 100.00 91.60 97.00 46.40 86.90 93.30 27.00
tic_tac_toe 7 164 206 57.10 28.60 100.00 45.10 25.00 86.40 45.10 27.20 50.00
bucket_sort 14 162 120 92.90 16.70 100.00 83.30 1.10 92.90 78.30 0.00 68.10
dynamic_stack 16 161 128 93.80 90.00 94.40 85.10 65.50 53.20 68.80 36.40 16.80
sudoku_solver 20 159 148 100.00 77.80 100.00 96.20 67.90 84.10 79.10 73.10 36.90
naval_battle 6 156 224 100.00 83.30 100.00 93.60 55.50 93.60 82.10 52.30 76.00
trie 19 152 128 100.00 100.00 59.10 99.30 94.60 48.50 88.30 90.90 35.20
Total/Average 151 1,900 1,572 94.38 67.43 92.41 86.30 51.59 72.92 80.00 47.34 47.59
Table 2. ATLAS Supported Language Constructs
Functions
Loops/ Conditionals
Struct/Enum/ Union
Func. Pointers
Goto/Label
Classes
Namespaces
Templates
Virtual Methods
Exceptions
Op. Overloading
C
C++
Figure 3. Performance metrics of ATLAS across 20 benchmark programs (avg. 119 LOC). Median execution time: 6.89s, peak memory: 135.01 MB. Trend lines show best-fit polynomial curves.
To evaluate the accuracy of ATLAS, we generate and verify manually the CFG and DFG for all the C and C++ projects from the TheAlgorithms dataset (TheAlgorithms, 2025b) (TheAlgorithms, 2025a). To evaluate the practicality of the tool, we compare the test generation performance of DeepSeek V3.2 with and without statement-level CFG as context. Table 2 summarizes the language constructs natively supported by ATLAS. We select the 10 largest C programs by LOC, excluding main functions and test scaffolding to isolate the core algorithmic logic. Additionally, we benchmark the tool’s scalability by monitoring its computational resource consumption relative to code complexity.
4.1. Impact of CodeView Combinations
Our analysis lies in advancing LLM-based control-flow reasoning for test generation, therefore our focus is on the comparison between state-of-the-art LLM with statement-level CFG (DS w/ CFG) and without CFG (DS w/o CFG). We use DeepSeek V3.2 with temperature=0, max_tokens=4096, and reasoning turned off. For DS w/ CFG, we extract the execution paths from the statement-level CFG and generate a test for each unique path. We replicate a similar architecture to PANTA with max_iterations=15 and LLM temperature=0.1 for test repairs. We discard any tests if they still fail after max iterations. For DS w/o CFG, we do not add the paths into the LLM’s context as we do not generate a CFG, and we keep the rest of the modules exactly identical.
We present the results in Table 1. In terms of line and branch coverages, DS w/ CFG consistently outperforms DS w/o CFG across all projects. On average, DS w/ CFG achieves 34.71% higher line coverage and 32.66% higher branch coverage than DS w/o CFG. Moreover, the performance is on-par with the industry-standard tool KLEE (Cadar et al., 2008) in most projects and even exceeds it in the larger projects: red_black_tree, multikey_quick_sort, hangman, and trie. Beyond coverage metrics, DS w/ CFG also demonstrates improved test efficiency: 75% of generated tests exercise unique execution paths compared to only 67% for DS w/o CFG. This indicates that CFG-guided test generation reduces redundant tests by enabling the LLM to target distinct control-flow paths systematically.
4.2. Scalability Analysis
To evaluate the scalability of ATLAS, we measure its performance during CFG generation on the 20 largest programs from the TheAlgorithms C dataset (TheAlgorithms, 2025b), ranging from 61 to 309 LOC. Figure 3 presents the execution time and memory consumption metrics across these programs. The median execution time is 6.89 seconds, with median peak memory usage of 135.01 MB. The trend lines in Figure 3 show best-fit polynomial curves, revealing that execution time and memory consumption scale polynomially with program size rather than exponentially, suggesting that ATLAS remains practical for moderately-sized programs typical of benchmark datasets.
5. Discussion and Limitations
ATLAS provides robust C/C++ support by prioritizing core program logic and inter-procedural dependencies. By requiring only syntactic validity rather than full compilability, the tool bypasses build-system overhead while maintaining broad industrial applicability. To preserve source-level context, ATLAS parses pre-processor directives without executing conditional compilation paths, ensuring it remains a flexible analyzer rather than a rigid compiler replacement. Furthermore, certain advanced C++ features, such as template metaprogramming, SFINAE, and complex inheritance hierarchies, are currently modeled with limited granularity to maintain graph clarity and computational scalability. These design choices ensure that ATLAS remains a performant infrastructure for generating the information-dense datasets required by modern ML4SE models.
6. Conclusion
We introduced ATLAS, a CLI that generates unified multi-view representations of C/C++ projects by aligning syntactic, control, and data-flow properties. By operating on raw source code without build-system dependencies, ATLAS enables the rapid generation of information-dense datasets for ML4SE. Our evaluation proves that ATLAS-guided LLMs significantly enhance control-flow reasoning, matching the effectiveness of symbolic execution tools like KLEE with polynomial scalability. ATLAS establishes a foundational infrastructure for graph-aware models, with future work targeting advanced C++ idioms and automated memory-safety analysis.
Appendix A ATLAS: Tool Demonstration Walkthrough
This appendix presents an evaluation of ATLAS’s success rates, compares its CFG representation with LLVM IR, and provides step-by-step demonstrations showcasing single-file analysis, combined multi-code view generation, and multi-file project analysis.
A.1. CFG and DFG Generation Success Rate
Table 3. Analysis of Failure Cases in ATLAS
Failure Category C C++
CFG DFG CFG DFG
goto statements 3 3 - -
Multi-threading 10 10 6 6
Pointer Arithmetic - 22 - 2
Operator Overloading - - 24 24
Static Variables - - - 2
Total 13 35 30 34
We evaluated ATLAS on the C and C++ subsets of The Algorithms datasets. The C dataset contains 406 files, where ATLAS achieved a success rate of 96.80% for CFG and 91.38% for DFG generation. The C++ dataset consists of 360 files, with success rates of 91.67% for CFGs and 90.56% for DFGs. The remaining uncovered cases are due to specific, complex language features that are not yet fully supported. Table 3 provides a detailed breakdown of the failure categories for both languages and analysis types.
A.2. CFG Representation Comparison: LLVM IR vs. ATLAS
Before demonstrating ATLAS’s capabilities, we illustrate the key difference between traditional compiler-based CFG generation and ATLAS’s statement-level approach. This comparison highlights why source-level CFGs are more suitable for program comprehension and software engineering tasks.
A.2.1. Input Source Code
We use a simple factorial program to compare CFG representations:
Figure 4. Source code of the factorial program used for CFG comparison.
A.2.2. LLVM IR-Based CFG (Clang)
Traditional compiler toolchains such as Clang generate CFGs at the LLVM Intermediate Representation (IR) level. While useful for compiler optimizations, LLVM IR CFGs operate on a lower-level abstraction that loses source-level code semantics:
Figure 5. LLVM IR-based CFG generated by Clang for the factorial function. Nodes contain low-level IR instructions (alloca, store, load, icmp, br) rather than source statements.
Limitations of LLVM IR CFGs:
  • Loss of source semantics — Original variable names are replaced with numbered registers (%1, %2, etc.)
  • Lowered abstractions — High-level constructs like for loops are decomposed into basic blocks with conditional branches
  • Memory operations — Simple variable assignments become sequences of alloca, store, and load instructions
  • Difficult traceability — Mapping IR nodes back to source code lines requires additional debug information
A.2.3. ATLAS Statement-Level CFG
In contrast, ATLAS generates CFGs at the source statement level, preserving the original code structure and semantics:
Figure 6. ATLAS-generated CFG for the same factorial program.
Advantages of ATLAS’s approach:
  • Source-level fidelity — Nodes contain actual source statements as written by developers
  • Preserved semantics — Loop structures (for), conditionals (if), and function calls remain explicit
  • Direct traceability — Each node includes line numbers for immediate source code correlation
  • Semantic edge labels — Control flow edges are annotated with meaningful labels (loop_update, function_return)
  • Inter-procedural visibility — Function calls and returns are explicitly represented as graph edges
This statement-level representation makes ATLAS’s CFGs directly applicable to software engineering tasks such as code review, bug localization, and program understanding, without requiring expertise in compiler internals.
A.3. Prerequisites
ATLAS can be run either via direct installation or Docker. For this demonstration, we use Docker for reproducibility:
docker build -t atlas .
All commands assume the working directory contains the example source files.
A.4. Scenario A: Single File Analysis — Control Flow Graph Generation
This scenario demonstrates ATLAS’s ability to generate a Control Flow Graph (CFG) from a single C source file, useful for understanding program execution paths.
A.4.1. Input Source Code
We use factorial.cpp, a simple factorial function with a loop and conditional return:
Figure 7. Source code of factorial.cpp.
A.4.2. Command
Terminal command
docker run --rm \
-v "\$(pwd):/work" -w /work \
atlas --lang cpp \
--code-file factorial.cpp \
--graphs cfg --output all
A.4.3. Output Files Generated
File Description
factorial.json Graph data in JSON format
factorial.dot GraphViz DOT representation
factorial.png Visual graph rendering
A.4.4. Generated Control Flow Graph
Figure 8. Control Flow Graph for factorial.cpp showing execution paths through the factorial function and main function. Nodes represent statements, edges represent control flow transitions. The loop structure in the factorial function and the conditional branch are clearly visible.
A.4.5. Interpretation
The CFG reveals:
  • Entry and exit points for each function (factorial and main)
  • Branching structure at the if (n <= 1) condition with two outgoing edges
  • Loop back-edge from the end of the for loop body back to the loop condition
  • Sequential flow through the main function’s variable declarations and function call
A.5. Scenario B: Combined Multi-Code View Generation
This scenario demonstrates ATLAS’s core capability: generating and combining multiple code representations (CFG and DFG) into a single unified graph. This combined view enables analyses that require correlating control flow and data dependency information simultaneously.
A.5.1. Input Source Code
We use a bubble sort implementation (bubble_sort.c) that demonstrates variable definitions, uses, and complex control flow:
Figure 9. Source code of bubble_sort.c.
A.5.2. Command
Terminal command
docker run --rm \
-v "\$(pwd):/work" -w /work \
atlas --lang c \
--code-file bubble_sort.c \
--graphs cfg,dfg --output all
A.5.3. Generated Combined CFG+DFG Graph
Figure 10. Combined Control Flow Graph and Data Flow Graph for bubble_sort.c rendered as a single unified graph.
A.5.4. Interpretation
The combined CFG+DFG visualization enables answering questions that require both control flow and data dependency information:
  • “Which variables are modified in the loop body?” — Follow red CFG edges to identify loop nodes, then trace blue DFG edges to find variable definitions
  • “Does a variable definition reach a specific use?” — Combine blue DFG definition-use edges with red CFG paths to verify reachability
  • “What data flows through the swap operation?” — Trace blue DFG edges showing how temp, arr[j], and arr[j+1] are connected through the swap sequence
A.6. Scenario C: Multi-File Project Analysis
This scenario demonstrates ATLAS’s ability to analyze entire projects by combining multiple source files into a unified analysis. This is essential for real-world codebases where functions span multiple files.
A.6.1. Input Project Structure
We analyze a small project with separate files for a math utilities library:
math_project/
|-- math_utils.h
|-- math_utils.c
|-- main.c
Figure 11. Header file math_utils.h declaring the add, multiply, and factorial function prototypes.
Figure 12. Implementation file math_utils.c containing arithmetic operations.
Figure 13. Main file main.c demonstrating usage of the math utilities API.
A.6.2. Command
Terminal command
docker run --rm \
-v "\$(pwd):/work" -w /work \
atlas --lang c \
--code-folder math_project \
--combined-name math_analysis \
--graphs cfg --output all
A.6.3. Output Files Generated
File Description
math_analysis.json Combined graph data from all source files
math_analysis.dot GraphViz DOT representation
math_analysis.png Visual graph rendering
A.6.4. Generated Combined Project CFG
Figure 14. Combined Control Flow Graph spanning all functions across math_utils.c and main.c. Each function (add, multiply, factorial, main) appears as a separate subgraph with its own entry and exit points. The graph reveals the control flow patterns including the loop structure in factorial and the sequential function calls in main.
A.6.5. Multi-File Analysis Benefits
Aspect Single-File Multi-File
Function relationships Limited to one file Cross-file visibility
Global variables Partial Complete
Header correlation Not possible Fully supported
Project-wide flow Not possible Enabled
A.6.6. Interpretation
The combined multi-file analysis enables:
  • Cross-file call tracking — Following calls from main.c to functions in math_utils.c
  • Complete function analysis — Viewing all functions and their control flow together
  • Inter-function dependencies — Tracing how factorial calls multiply
  • API usage patterns — Correlating header declarations with implementations

### v3

@@
 Traditional static analysis tools often fail in these scenarios due to their reliance on successful compilation, while Large Language Models (LLMs) lack the structural context necessary to reason about complex program logic.
-We introduce ATLAS, a Python-based Command-Line Interface (CLI).
-It is characterized by: (i) inter-procedural, type-aware analysis that captures dependencies across function boundaries; (ii) support for both full-project and partial analysis of non-compilable projects; (iii) graph optimizations, such as variable collapsing and node blacklisting; and (iv) synchronized multi-view graphs that align syntax, execution paths, and data-flow logic.
-Evaluating ATLAS with DeepSeek V3.2 for automated test generation demonstrates a 34.71% increase in line coverage and 32.66% in branch coverage, achieving performance on par with and even exceeding the industry-standard symbolic execution tool KLEE on complex projects.
-With polynomial scalability, ATLAS provides a robust infrastructure for generating the information-dense datasets required by next-generation, graph-aware Machine Learning for Software Engineering (ML4SE) models.
-Video demonstration: https://youtu.be/xEJDNYkiD4s
-Tool github repository: https://github.com/jaid-monwar/ATLAS-multi-view-code-representation-tool.git
+We introduce ATLAS, a Python-based CLI that generates unified multi-view representations for large-scale C/C++ projects with high accuracy, achieving success rates up to 96.80% for CFGs and 91.38% for DFGs.
+ATLAS is characterized by: (i) inter-procedural, type-aware analysis across function boundaries; (ii) support for both full and partial analysis of non-compilable projects; (iii) graph optimizations such as variable collapsing and node blacklisting; and (iv) synchronized multi-view graphs that align syntax, execution paths, and data-flow logic.
+Evaluating ATLAS with DeepSeek V3.2 for automated test generation demonstrates a 34.71% increase in line coverage and 32.66% in branch coverage, matching or exceeding the performance of the symbolic execution tool KLEE on complex projects.
+With polynomial scalability, ATLAS provides a robust infrastructure for generating the information-dense datasets required by next-generation, graph-aware ML4SE models.
+Video demonstration: https://youtu.be/QGuJZhj9CTA
+Tool Github repository: ATLAS
 Static Program Analysis, Control Flow, Data Flow Analysis, C/C++
-conference: ; ;
+copyright: noneconference: ; ;
 1. Introduction
-Programming languages like C and C++ are considered still as the foundation of performance-critical software, such as operating systems and embedded firmware (Stroustrup, 2020), where hardware proximity and resource efficiency are paramount.
-However, as C/C++ based systems scale, managing their evolution becomes increasingly complex (Brooks, 1987), particularly when analyzing sub-modules that are non-compilable or contain unresolved dependencies.
-Consequently, the field has transitioned toward the Machine Learning for Software Engineering (ML4SE) to automate tasks like test generation (Fontes and Gay, 2023) (Durelli et al., 2019), code translation (Tufano et al., 2019) (Ibrahimzada et al., 2025), source code classification (Mou et al., 2016), and code summarization (Haiduc et al., 2010) (Jiang et al., 2017).
+Programming languages like C and C++ are still considered the foundation of performance-critical software, such as operating systems and embedded firmware (Stroustrup, 2020), where hardware proximity and resource efficiency are paramount.
+However, as C/C++ based systems scale, managing their evolution becomes increasingly complex (Brooks, 1987), particularly when analyzing sub-modules that contain unresolved dependencies.
+Consequently, the field has transitioned towards Machine Learning for Software Engineering (ML4SE) to automate tasks like test generation (Fontes and Gay, 2023) (Durelli et al., 2019), code translation (Tufano et al., 2019) (Ibrahimzada et al., 2025), source code classification (Mou et al., 2016), and code summarization (Haiduc et al., 2010) (Jiang et al., 2017).
 Within this paradigm, Large Language Models (LLMs) offer a flexible alternative to traditional tools.
-These can derive semantic context even from unbuildable projects by processing code as structured textual sequences (Wang et al., 2025) (Panichella et al., 2013) (Li et al., 2025).
+LLMs can derive semantic context even from unbuildable projects by processing code as structured textual sequences (Wang et al., 2025) (Panichella et al., 2013) (Li et al., 2025).
 However, LLMs are limited by their probabilistic nature, often prioritizing token adjacency over the deep, non-linear relationships that dictate program behavior (Anand et al., 2024) (Pendyala and Thakur, 2025).
 Bridging this ’semantic gap’ requires effective representation learning.
-Such techniques are essential for training and evaluating ML models while also enhancing the reasoning capabilities of LLMs.
+Such techniques are essential for training and evaluating Machine Learning (ML) models while also enhancing the reasoning capabilities of LLMs.
 Empirical evidence (Liu and Jabbarvand, 2025) (Samoaa et al., 2022) shows that performance improves significantly when models utilize multi-view data that explicitly encodes these underlying logical relationships.
-To bridge this gap, research has shifted toward representing code as integrated graphs (Chimalakonda et al., 2023).
+Consequently, research has shifted toward representing code as integrated graphs (Chimalakonda et al., 2023).
 The Abstract Syntax Tree (AST) encodes the hierarchical syntactic structure of code constructs, the Control Flow Graph (CFG) models execution paths, and the Data Flow Graph (DFG) captures the variable dependencies.
-However, their integration into a multi-view representation allows models to holistically capture program logic.
+Crucially, their integration into a multi-view representation allows models to holistically capture program logic.
 This approach is validated by CodeSAM (Mathai et al., 2024), which uses attention-masking to improve code search, and PANTA (Gu et al., 2025), which leverages CFG analysis to increase the test coverage by 23%.
-However, these successes are largely centered on Java, and C/C++ projects lack such high-level abstractions and require significantly more code to implement equivalent logic, exhausting LLM context windows more rapidly(Chimalakonda et al., 2023).
-Moreover, this difficulty is further complicated by C-specific hurdles like pointer aliasing, and type obscuration, which remain under-addressed (Schubert et al., 2021) in representation learning research.
+However, these successes are largely centered on Java.
+C/C++ projects, on the other hand, lack such high-level abstractions and require significantly more code to implement equivalent logic, exhausting LLM context windows more rapidly (Chimalakonda et al., 2023).
+Moreover, this difficulty is further complicated by C-specific hurdles like pointer aliasing and type obscuration, which remain under-addressed (Schubert et al., 2021) in representation learning research.
 To address these challenges, this paper introduces ATLAS, a Python-based Command-Line Interface (CLI) designed to function as a high-performance data generation engine for ML4SE.
@@
 In contrast, ATLAS operates on source ASTs, preserving semantic metadata for non-compilable projects and natively modeling constructs like templates and lambdas.
-Additionally, static tools like Joern (14) provide multi-language code-views but lack a native mechanism to customize or combine these representations into a unified structure.
+Additionally, static tools like Joern (joe, [n.
+d.]) provide multi-language code-views but lack a native mechanism to customize or combine these representations into a unified structure.
 This often results in isolated, function-level graphs that cannot resolve the long-range dependencies required for a comprehensive project-level analysis.
@@
 First, it provides statement-level DFGs for C and C++, whereas existing research tools like PhASAR (Schubert et al., 2019) focus primarily on IR.
-Second, it resolves the fragmentation of current toolchains, where developers must currently stitch together disparate tools, such as Clang for ASTs and Joern (14) for CFGs, using custom scripts.
+Second, it resolves the fragmentation of current toolchains, where developers must stitch together disparate tools, such as Clang for ASTs and Joern (joe, [n.
+d.]) for CFGs, using custom scripts.
 This fragmented process is inefficient and lacks cross-view alignment.
-ATLAS provides a unified framework for aligned multi-view graphs, eliminating redundant analysis and streamlining ML4SE dataset pipelines.
+ATLAS provides a unified framework for aligned multi-view graphs, eliminating redundant analysis and streamlining ML4SE tasks.
 3. ATLAS Tool
@@
 (c) DFG depicting the flow of data values and dependencies between variables and operations.
-ATLAS follows a two-step workflow (Fig.
+ATLAS follows a three-step workflow (Fig.
 1).
-At first, the Preprocessor and Parser normalize multi-file projects into a Parsed Code Model enriched with scope, type, and declaration-use metadata.
+First, the Preprocessor and Parser normalize multi-file projects into a Parsed Code Model enriched with scope, type, and declaration-use metadata.
 Second, the Combined Code View Generator constructs aligned graphs via four modules: the AST, CFG, and DFG Drivers, and the Unified Multi-view Graph.
-These components utilize Reaching Definition Analysis (RDA) and inter-procedural mapping to ensure consistent node alignment across all views.
 Finally, ATLAS exports these representations in JSON, DOT, or PNG, linking a program’s syntactic structure directly to its execution behavior.
 3.1. Preprocessor and Parser
-The Preprocessor and Parser generate a unified Parsed Code Model.
+The module generates a unified Parsed Code Model.
 The Preprocessor normalizes the codebase by consolidating multi-file projects, removing non-semantic elements (e.g., comments and empty lines), and resolving global macro and type definitions.
 Subsequently, the Parser leverages Tree-sitter to generate concrete syntax trees (CSTs), which are traversed to build an enriched symbol table mapping scope hierarchies, data types, and declaration-use relationships.
-This metadata layer serves as the foundational reference for all subsequent graph generation.
 3.2. Code Representation Generation
 The Combined Code View Generator utilizes the enriched symbol table to construct aligned AST, CFG, and DFG views.
-This shared reference ensures node-level consistency for unified multi-view analysis.
+These components utilize Reaching Definition Analysis (RDA) and inter-procedural mapping to ensure consistent node alignment across all views.
 Construction logic and refinements for each representation are as follows.
 3.2.1. AST
-The CST produced in the preprocessing step preserves all syntactic details, which ATLAS traverses to build an AST index mapping each node’s position and type to a unique identifier.
-To refine this representation, ATLAS prunes non-semantic tokens, such as brackets and semicolons, to produce a core AST, as shown in Fig.
-2(a).
-The AST Driver further provides two customization mechanisms to control graph granularity: (i) Variable Collapsing, which unifies all occurrences of a specific variable identifier into a single node to reduce redundancy, and (ii) Node Blacklisting, which allows users to omit irrelevant node categories (e.g., comment, modifier) and their children.
+The CST produced in the preprocessing step preserves all syntactic details, which ATLAS traverses to build an AST index, mapping each node’s position and type to a unique identifier.
+To refine this representation, ATLAS prunes non-semantic tokens, such as brackets and semicolons, to produce a core AST (Fig.
+2(a)).
+The AST Driver further provides two customization mechanisms to control graph granularity: (i) Variable Collapsing, which unifies all occurrences of a specific variable identifier into a single node to reduce redundancy, and (ii) Node Blacklisting, which allows users to omit irrelevant node categories (e.g., modifier) and their children.
 These refinements allow users to distill the program into its core logical elements, effectively mitigating the computational overhead and overfitting risks associated with high-dimensional syntax.
 3.2.2. CFG
-The CFG Driver generates statement-level, inter-procedural representations where each node corresponds to a source statement and edges map all potential execution transitions.
+The CFG Driver generates statement-level and inter-procedural representations where each node corresponds to a source statement and edges map all potential execution transitions.
 Unlike traditional tools that produce isolated function-level graphs, ATLAS establishes global connectivity by resolving function calls as directed edges from caller statements to callee definitions.
 This capability is demonstrated in Fig.
-2(b), where the ternary operator on line 20 triggers indirect calls to f1() and f2() via function pointers; ATLAS correctly models this ambiguity by generating branching edges to both function definitions.
+2(b), where the ternary operator on line 20 triggers indirect calls to f1() and f2() via function pointers.
+ATLAS correctly models this ambiguity by generating branching edges to both function definitions.
 Furthermore, the tool natively supports recursion by representing it as self-referential edges, as seen in the recursive call within f1() on line 7.
@@
 3.2.4. Combined Multi-View Graph
-ATLAS merges these complementary views into a unified multi-view graph by preserving consistent node identities across all representations.
-This alignment allows downstream ML/DL models to learn from syntactic, control-flow, and data-flow properties simultaneously.
-The tool supports fifteen combination variants, including pairwise and triple-view integrations, all of which remain compatible with user-defined collapsing and blacklisting.
-By assigning distinct type attributes to each edge class, ATLAS enables models to selectively prioritize specific program semantics over detail-heavy syntax.
+ATLAS merges AST, CFG, and DFG views into a unified multi-view graph by preserving consistent node identities across all representations.
+This alignment enables downstream models to learn syntactic, control-flow, and data-flow properties simultaneously.
+The tool supports fifteen combination variants, including pairwise and triple-view integrations, fully compatible with user-defined collapsing and blacklisting.
+By assigning distinct type attributes to each edge class, ATLAS allows models to prioritize specific program semantics over detail-heavy syntax.
 3.3. Visualizer
 The Visualizer manages the exportation of aligned graphs into three different formats.
-While the JSON output is optimized for Machine Learning (ML) integration, the DOT and PNG formats facilitate human-centric debugging and structural inspection.
+While the JSON output is optimized for ML integration, the DOT and PNG formats facilitate human-centric debugging and structural inspection.
 For multi-view representations, the component employs color-coded edge classes to distinguish between syntactic, control, and data-flow properties.
-Additionally, the Visualizer implements specialized C++ formatting, such as escaping namespace qualifiers and reserved keywords, to ensure high-fidelity rendering of complex codebase structures.
 4. ATLAS In Practice
@@
 Total/Average 151 1,900 1,572 94.38 67.43 92.41 86.30 51.59 72.92 80.00 47.34 47.59
-Table 2. ATLAS Supported Language Constructs
-Functions
-Loops/ Conditionals
-Struct/Enum/ Union
-Func. Pointers
-Goto/Label
-Classes
-Namespaces
-Templates
-Virtual Methods
-Exceptions
-Op. Overloading
-C
-C++
 Figure 3. Performance metrics of ATLAS across 20 benchmark programs (avg. 119 LOC). Median execution time: 6.89s, peak memory: 135.01 MB. Trend lines show best-fit polynomial curves.
-To evaluate the accuracy of ATLAS, we generate and verify manually the CFG and DFG for all the C and C++ projects from the TheAlgorithms dataset (TheAlgorithms, 2025b) (TheAlgorithms, 2025a).
-To evaluate the practicality of the tool, we compare the test generation performance of DeepSeek V3.2 with and without statement-level CFG as context.
-Table 2 summarizes the language constructs natively supported by ATLAS.
+To evaluate the accuracy of ATLAS, we generate and manually verify the CFG and DFG for all the C and C++ projects from the TheAlgorithms dataset (TheAlgorithms, 2025) (TheAlgorithms, 2025).
+To evaluate the practicality of the tool, we compare the test generation performance of DeepSeek V3.2 (DeepSeek-AI et al., 2025) with and without statement-level CFG as context.
 We select the 10 largest C programs by LOC, excluding main functions and test scaffolding to isolate the core algorithmic logic.
 Additionally, we benchmark the tool’s scalability by monitoring its computational resource consumption relative to code complexity.
-4.1. Impact of CodeView Combinations
-Our analysis lies in advancing LLM-based control-flow reasoning for test generation, therefore our focus is on the comparison between state-of-the-art LLM with statement-level CFG (DS w/ CFG) and without CFG (DS w/o CFG).
-We use DeepSeek V3.2 with temperature=0, max_tokens=4096, and reasoning turned off.
-For DS w/ CFG, we extract the execution paths from the statement-level CFG and generate a test for each unique path.
-We replicate a similar architecture to PANTA with max_iterations=15 and LLM temperature=0.1 for test repairs.
-We discard any tests if they still fail after max iterations.
-For DS w/o CFG, we do not add the paths into the LLM’s context as we do not generate a CFG, and we keep the rest of the modules exactly identical.
-We present the results in Table 1.
-In terms of line and branch coverages, DS w/ CFG consistently outperforms DS w/o CFG across all projects.
-On average, DS w/ CFG achieves 34.71% higher line coverage and 32.66% higher branch coverage than DS w/o CFG.
-Moreover, the performance is on-par with the industry-standard tool KLEE (Cadar et al., 2008) in most projects and even exceeds it in the larger projects: red_black_tree, multikey_quick_sort, hangman, and trie.
-Beyond coverage metrics, DS w/ CFG also demonstrates improved test efficiency: 75% of generated tests exercise unique execution paths compared to only 67% for DS w/o CFG.
-This indicates that CFG-guided test generation reduces redundant tests by enabling the LLM to target distinct control-flow paths systematically.
-4.2. Scalability Analysis
-To evaluate the scalability of ATLAS, we measure its performance during CFG generation on the 20 largest programs from the TheAlgorithms C dataset (TheAlgorithms, 2025b), ranging from 61 to 309 LOC.
-Figure 3 presents the execution time and memory consumption metrics across these programs.
-The median execution time is 6.89 seconds, with median peak memory usage of 135.01 MB.
-The trend lines in Figure 3 show best-fit polynomial curves, revealing that execution time and memory consumption scale polynomially with program size rather than exponentially, suggesting that ATLAS remains practical for moderately-sized programs typical of benchmark datasets.
-5. Discussion and Limitations
-ATLAS provides robust C/C++ support by prioritizing core program logic and inter-procedural dependencies.
-By requiring only syntactic validity rather than full compilability, the tool bypasses build-system overhead while maintaining broad industrial applicability.
-To preserve source-level context, ATLAS parses pre-processor directives without executing conditional compilation paths, ensuring it remains a flexible analyzer rather than a rigid compiler replacement.
-Furthermore, certain advanced C++ features, such as template metaprogramming, SFINAE, and complex inheritance hierarchies, are currently modeled with limited granularity to maintain graph clarity and computational scalability.
-These design choices ensure that ATLAS remains a performant infrastructure for generating the information-dense datasets required by modern ML4SE models.
-6. Conclusion
-We introduced ATLAS, a CLI that generates unified multi-view representations of C/C++ projects by aligning syntactic, control, and data-flow properties.
-By operating on raw source code without build-system dependencies, ATLAS enables the rapid generation of information-dense datasets for ML4SE.
-Our evaluation proves that ATLAS-guided LLMs significantly enhance control-flow reasoning, matching the effectiveness of symbolic execution tools like KLEE with polynomial scalability.
-ATLAS establishes a foundational infrastructure for graph-aware models, with future work targeting advanced C++ idioms and automated memory-safety analysis.
-Appendix A ATLAS: Tool Demonstration Walkthrough
-This appendix presents an evaluation of ATLAS’s success rates, compares its CFG representation with LLVM IR, and provides step-by-step demonstrations showcasing single-file analysis, combined multi-code view generation, and multi-file project analysis.
-A.1. CFG and DFG Generation Success Rate
+4.1. Graph Generation Success Rates
 Table 2. Analysis of Failure Cases in ATLAS
@@
 Total 13 35 30 34
-We evaluated ATLAS on the C and C++ subsets of The Algorithms datasets.
-The C dataset contains 406 files, where ATLAS achieved a success rate of 96.80% for CFG and 91.38% for DFG generation.
-The C++ dataset consists of 360 files, with success rates of 91.67% for CFGs and 90.56% for DFGs.
-The remaining uncovered cases are due to specific, complex language features that are not yet fully supported.
-Table 3 provides a detailed breakdown of the failure categories for both languages and analysis types.
-A.2. CFG Representation Comparison: LLVM IR vs. ATLAS
+Evaluating ATLAS on the TheAlgorithms dataset (406 C and 360 C++ files) yielded high success rates for graph generation: CFG: 96.80% (C) and 91.67% (C++); DFG: 91.38% (C) and 90.56% (C++).
+As detailed in Table 2, remaining failures stem from the unmanaged memory model of C/C++, where runtime behaviors, such as dynamic pointer arithmetic and array element modifications, cannot be fully resolved statically.
+Additionally, to prioritize graph clarity and computational scalability, certain complex constructs like goto statements and static class variables are currently modeled with limited granularity.
+By requiring only syntactic validity rather than a resolved build environment, ATLAS bypasses the overhead of compilation while maintaining broad industrial applicability.
+These design trade-offs establish ATLAS as a performant tool for generating the information-dense datasets essential for ML models.
+4.2. Impact of CodeView Combinations
+Our analysis lies in advancing LLM-based control-flow reasoning for test generation; therefore, our focus is on the comparison between state-of-the-art LLM with statement-level CFG (DS w/ CFG) and without CFG (DS w/o CFG).
+We use DeepSeek V3.2 with temperature=0, max_tokens=4096, and reasoning turned off.
+For DS w/ CFG, we extract the execution paths from the statement-level CFG and generate a test for each unique path.
+We replicate a similar architecture to PANTA (Gu et al., 2025) with max_iterations=15 and LLM temperature=0.1 for test repairs.
+We discard any tests if they still fail after max iterations.
+For DS w/o CFG, we do not add the paths into the LLM’s context as we do not generate a CFG, and we keep the rest of the modules exactly identical.
+We present the results in Table 1.
+In terms of line and branch coverages, DS w/ CFG consistently outperforms DS w/o CFG almost across all projects.
+On average, DS w/ CFG achieves 34.71% higher line coverage and 32.66% higher branch coverage than DS w/o CFG.
+The coverage performance of DS w/ CFG is on-par with the industry-standard tool KLEE (Cadar et al., 2008) in most projects and even exceeds it in the larger projects (red_black_tree, hangman, etc).
+Beyond coverage metrics, DS w/ CFG also demonstrates improved test efficiency of around 75% in terms of the ratio of tests that follow at least one unique execution path to the total number of generated tests.
+This is in stark contrast to only 67% for DS w/o CFG.
+This indicates that ATLAS-generated statement-level CFG-guided test generation reduces redundant tests by enabling the LLM to target distinct control-flow paths systematically.
+4.3. Scalability Analysis
+To evaluate the scalability of ATLAS, we measure its performance during CFG generation on the 20 largest C programs from the TheAlgorithms C dataset (TheAlgorithms, 2025), ranging from 61 to 309 LOC.
+Figure 3 presents the execution time and memory consumption metrics across these programs.
+The median execution time is 6.89 seconds, with median peak memory usage of 135.01 MB.
+The trend lines in Figure 3 show best-fit polynomial curves, revealing that execution time and memory consumption scale polynomially with program size rather than exponentially, suggesting that ATLAS remains practical for moderately-sized programs typical of benchmark datasets.
+5. Conclusion
+We introduced ATLAS, a Python-based CLI that generates unified multi-view representations of C/C++ projects by aligning syntactic, control, and data-flow properties.
+By operating on raw source code without build-system dependencies, ATLAS enables the rapid generation of information-dense datasets for ML4SE.
+Our evaluation demonstrates that ATLAS-guided LLMs significantly enhance control-flow reasoning, matching the effectiveness of symbolic execution tools like KLEE, while maintaining polynomial scalability.
+ATLAS establishes a foundational infrastructure for graph-aware models, with future work targeting advanced C++ idioms and automated memory-safety analysis.
+Appendix A ATLAS: Tool Demonstration Walkthrough
+This appendix compares ATLAS’s CFG representation with LLVM IR and provides step-by-step demonstrations showcasing single-file analysis, combined multi-code view generation, and multi-file project analysis.
+Video demonstration: https://youtu.be/QGuJZhj9CTA
+Tool Github repository: ATLAS
+A.1. CFG Representation Comparison: LLVM IR vs. ATLAS
 Before demonstrating ATLAS’s capabilities, we illustrate the key difference between traditional compiler-based CFG generation and ATLAS’s statement-level approach.
 This comparison highlights why source-level CFGs are more suitable for program comprehension and software engineering tasks.
-A.2.1. Input Source Code
+A.1.1. Input Source Code
 We use a simple factorial program to compare CFG representations:
 Figure 4. Source code of the factorial program used for CFG comparison.
-A.2.2. LLVM IR-Based CFG (Clang)
+A.1.2. LLVM IR-Based CFG (Clang)
 Traditional compiler toolchains such as Clang generate CFGs at the LLVM Intermediate Representation (IR) level.
@@
   • Difficult traceability — Mapping IR nodes back to source code lines requires additional debug information
-A.2.3. ATLAS Statement-Level CFG
+A.1.3. ATLAS Statement-Level CFG
 In contrast, ATLAS generates CFGs at the source statement level, preserving the original code structure and semantics:
@@
 This statement-level representation makes ATLAS’s CFGs directly applicable to software engineering tasks such as code review, bug localization, and program understanding, without requiring expertise in compiler internals.
-A.3. Prerequisites
+A.2. Prerequisites
 ATLAS can be run either via direct installation or Docker. For this demonstration, we use Docker for reproducibility:
@@
 All commands assume the working directory contains the example source files.
-A.4. Scenario A: Single File Analysis — Control Flow Graph Generation
+A.3. Scenario A: Single File Analysis — Control Flow Graph Generation
 This scenario demonstrates ATLAS’s ability to generate a Control Flow Graph (CFG) from a single C source file, useful for understanding program execution paths.
-A.4.1. Input Source Code
-We use factorial.cpp, a simple factorial function with a loop and conditional return:
+A.3.1. Input Source Code
+We use factorial.cpp, a file containing a simple factorial function with a loop and conditional return:
 Figure 7. Source code of factorial.cpp.
-A.4.2. Command
-Terminal command
+A.3.2. Command
+Run the following terminal command:
 docker run --rm \
@@
 --graphs cfg --output all
-A.4.3. Output Files Generated
+A.3.3. Output Files Generated
 File Description
@@
 factorial.png Visual graph rendering
-A.4.4. Generated Control Flow Graph
+A.3.4. Generated Control Flow Graph
 Figure 8.
@@
 The loop structure in the factorial function and the conditional branch are clearly visible.
-A.4.5. Interpretation
+A.3.5. Interpretation
 The CFG reveals:
@@
   • Sequential flow through the main function’s variable declarations and function call
-A.5. Scenario B: Combined Multi-Code View Generation
+A.4. Scenario B: Combined Multi-Code View Generation
 This scenario demonstrates ATLAS’s core capability: generating and combining multiple code representations (CFG and DFG) into a single unified graph.
 This combined view enables analyses that require correlating control flow and data dependency information simultaneously.
-A.5.1. Input Source Code
+A.4.1. Input Source Code
 We use a bubble sort implementation (bubble_sort.c) that demonstrates variable definitions, uses, and complex control flow:
 Figure 9. Source code of bubble_sort.c.
-A.5.2. Command
-Terminal command
+A.4.2. Command
+Run the following terminal command:
 docker run --rm \
@@
 --graphs cfg,dfg --output all
-A.5.3. Generated Combined CFG+DFG Graph
+A.4.3. Generated Combined CFG+DFG Graph
 Figure 10. Combined Control Flow Graph and Data Flow Graph for bubble_sort.c rendered as a single unified graph.
-A.5.4. Interpretation
+A.4.4. Interpretation
 The combined CFG+DFG visualization enables answering questions that require both control flow and data dependency information:
@@
   • “What data flows through the swap operation?” — Trace blue DFG edges showing how temp, arr[j], and arr[j+1] are connected through the swap sequence
-A.6. Scenario C: Multi-File Project Analysis
+A.5. Scenario C: Multi-File Project Analysis
 This scenario demonstrates ATLAS’s ability to analyze entire projects by combining multiple source files into a unified analysis.
 This is essential for real-world codebases where functions span multiple files.
-A.6.1. Input Project Structure
+A.5.1. Input Project Structure
 We analyze a small project with separate files for a math utilities library:
@@
 Figure 13. Main file main.c demonstrating usage of the math utilities API.
-A.6.2. Command
-Terminal command
+A.5.2. Command
+Run the following terminal command:
 docker run --rm \
@@
 --graphs cfg --output all
-A.6.3. Output Files Generated
+A.5.3. Output Files Generated
 File Description
@@
 math_analysis.png Visual graph rendering
-A.6.4. Generated Combined Project CFG
+A.5.4. Generated Combined Project CFG
 Figure 14.
@@
 The graph reveals the control flow patterns including the loop structure in factorial and the sequential function calls in main.
-A.6.5. Multi-File Analysis Benefits
+A.5.5. Multi-File Analysis Benefits
 Aspect Single-File Multi-File
@@
 Project-wide flow Not possible Enabled
-A.6.6. Interpretation
+A.5.6. Interpretation
 The combined multi-file analysis enables: